# CTC-DRO: Robust Optimization for Reducing Language Disparities in Speech Recognition

### v1

@@ Abstract
Modern deep learning models often achieve high overall performance, but consistently fail on specific subgroups. Group distributionally robust optimization (group DRO) addresses this problem by minimizing the worst-group loss, but it fails when group losses misrepresent performance differences between groups. This is common in domains like speech, where the widely used connectionist temporal classification (CTC) loss scales with input length and varies with linguistic and acoustic properties, leading to spurious differences between group losses. We present CTC-DRO, which addresses the shortcomings of the group DRO objective by smoothing the group weight update to prevent overemphasis on consistently high-loss groups, while using input length-matched batching to mitigate CTC’s scaling issues. We evaluate CTC-DRO on the task of multilingual automatic speech recognition (ASR) across five language sets from the ML-SUPERB 2.0 benchmark. CTC-DRO consistently outperforms group DRO and CTC-based baseline models, reducing the worst-language error by up to $65.9\%$ and the average error by up to $47.7\%$. CTC-DRO can be applied to ASR with minimal computational costs, and offers the potential for reducing group disparities in other domains with similar challenges.
@@ 1 Introduction
To address these issues, we present CTC-DRO, which optimizes a generalization of the group DRO objective, specifically by smoothing the up-weighting of high-loss groups. This new objective prevents overemphasis on groups with consistently and disproportionately high training losses. Also, we use length-matched group losses to mitigate the scaling properties of CTC. We evaluate CTC-DRO using language subsets randomly selected from the ML-SUPERB 2.0 (Shi et al., 2024) benchmark. In this setting, CTC-DRO models outperform both group DRO and CTC-based baseline models across five language sets. Specifically, CTC-DRO models reduce the error rate of the worst-performing language in all of the five sets, with improvements of up to $65.9\%$, while also improving the average performance across all languages by up to $47.7\%$. Our code and newly trained models will be made publicly available.
@@ 5 Results
1 MMS Baseline 59.7 (nan) 24.1 96.6
group DRO 86.0 (nan) 31.9 62.8
CTC-DRO 56.2 (nan) 21.9 97.6
XLS-R Baseline 187.9 (cmn) 58.5 77.8
group DRO 86.6 (cmn) 40.3 86.8
CTC-DRO 64.0 (nan) 30.6 87.8
2 MMS Baseline 45.2 (yue) 15.5 98.4
group DRO 56.9 (yue) 21.9 97.3
CTC-DRO 43.5 (yue) 14.8 98.2
XLS-R Baseline 64.8 (yue) 17.4 96.6
group DRO 60.0 (yue) 22.2 88.6
CTC-DRO 43.7 (yue) 15.4 96.4
3 MMS Baseline 36.6 (kor) 16.8 99.2
group DRO 33.4 (kor) 20.9 98.8
CTC-DRO 32.8 (khm) 16.0 99.2
XLS-R Baseline 34.0 (khm) 16.4 97.9
group DRO 36.9 (khm) 24.1 96.5
CTC-DRO 32.4 (kor) 16.6 96.5
4 MMS Baseline 28.4 (urd) 15.2 89.9
group DRO 30.5 (urd) 15.6 92.1
CTC-DRO 26.3 (urd) 14.6 94.7
@@
5 MMS Baseline 98.2 (jpn) 28.6 98.4
group DRO 61.1 (jpn) 29.9 67.0
CTC-DRO 53.0 (jpn) 21.9 98.9
XLS-R Baseline 120.9 (jpn) 32.5 89.0
group DRO 98.1 (jpn) 38.1 57.7
CTC-DRO 77.1 (jpn) 26.0 91.0
@@
For each language set, CTC-DRO models achieve a lower CER for the worst-performing language compared to the baseline and group DRO models. The largest improvement is obtained on set 1 using XLS-R in the balanced data setting, showing a relative CER reduction of $65.9\%$ compared to the baseline model. Note that CTC-DRO also results in the best average CER in 13 out of 14 settings (seven sets with two models each) compared to both the baseline and group DRO models, leading to relative CER reductions up to $47.7\%$. The only exception is for XLS-R in balanced set 3, where the average CER is slightly worse with CTC-DRO (16.6%) than the baseline (16.4%). In terms of LID accuracy, CTC-DRO models improve over the baseline models in 10 out of 14 settings, with a maximum improvement of 10 points for XLS-R in balanced set 1. In the remaining settings, the LID accuracy already exceeds 90%, leaving little room for further improvement with CTC-DRO.
In contrast, group DRO worsens the CER of the worst-performing language in seven out of 14 settings compared to the baseline model, with the highest relative CER increase of $58\%$ on set 2 using MMS trained on all available training data. Also, group DRO increases the average CER compared to the baseline in 13 out of 14 settings. This finding shows the ineffectiveness of the original group DRO formulation in this challenging setting, and the substantial added robustness of the modifications in CTC-DRO.
In four settings, the worst-performing language changes between the baseline and CTC-DRO models. For example, in set 3 with MMS trained on balanced data, it shifts from Korean to Khmer. As shown in Table 7, the CTC-DRO model reduces the CER for Korean from 36.6 to 29.6, while slightly increasing the CER for Khmer from 32.1 to 32.8. Overall, CTC-DRO consistently improves the performance of the worst-performing language, occasionally at a small cost to the performance of other languages, while still achieving a lower CER on average (see Appendix B).
@@ 6.1 Ablation Study
We find that removing either component from CTC-DRO leads to a substantial decrease in performance (see Table 4, detailed results in Appendix B). Specifically, the CER of the worst-performing language increases by up to $151.9\%$ and the average CER by up to $136.2\%$ compared to a model trained using the complete CTC-DRO algorithm. The ablation study also shows that the smoothed maximization objective has the strongest effect on reducing both the CER of the worst-performing language and the average CER.
@@ Appendix B Results
1 MMS Baseline 8.4 56.7 59.7 3.7 14.3 1.8 24.1 96.6
group DRO 24.5 48.1 86.0 5.4 18.1 9.1 31.9 62.8
CTC-DRO 10.4 45.1 56.2 3.6 14.1 1.9 21.9 97.6
XLS-R Baseline 26.6 187.9 84.1 11.2 30.2 11.2 58.5 77.8
group DRO 27.4 86.6 82.7 11.6 25.5 7.8 40.3 86.8
CTC-DRO 18.4 59.7 64.0 7.9 26.6 7.2 30.6 87.8
@@
2 MMS Baseline 0.2 22.0 10.6 4.6 10.6 45.2 15.5 98.4
group DRO 10.8 29.9 12.4 8.9 12.4 56.9 21.9 97.3
CTC-DRO 0.8 22.7 9.0 6.7 5.8 43.5 14.8 98.2
XLS-R Baseline 0.4 19.0 6.8 4.7 8.9 64.8 17.4 96.6
group DRO 11.9 31.4 12.3 5.7 11.7 60.0 22.2 88.6
CTC-DRO 0.8 21.8 11.9 5.8 8.3 43.7 15.4 96.4
@@
3 MMS Baseline 32.1 11.6 36.6 8.1 2.4 9.9 16.8 99.2
group DRO 30.7 18.9 33.4 18.5 10.0 13.5 20.9 98.8
CTC-DRO 32.8 11.8 29.6 8.5 2.9 10.2 16.0 99.2
XLS-R Baseline 34.0 11.4 32.6 8.0 2.2 10.4 16.4 97.9
group DRO 36.9 21.4 35.3 24.2 10.5 16.6 24.1 96.5
CTC-DRO 31.4 11.9 32.4 8.4 3.0 12.6 16.6 96.5
@@
5 MMS Baseline 6.9 11.8 33.7 98.2 12.7 7.9 28.6 98.4
group DRO 28.7 27.1 35.1 61.1 17.7 9.5 29.9 67.0
CTC-DRO 10.0 14.1 31.9 53.0 13.9 8.7 21.9 98.9
XLS-R Baseline 5.2 11.4 38.0 120.9 11.8 7.9 32.5 89.0
group DRO 29.2 27.4 44.7 98.1 17.7 11.2 38.1 57.7
CTC-DRO 6.1 11.2 41.5 77.1 11.1 8.9 26.0 91.0

### v2

@@ Abstract
Modern deep learning models often achieve high overall performance, but consistently fail on specific subgroups. Group distributionally robust optimization (group DRO) addresses this problem by minimizing the worst-group loss, but it fails when group losses misrepresent performance differences between groups. This is common in domains like speech, where the widely used connectionist temporal classification (CTC) loss scales with input length and varies with linguistic and acoustic properties, leading to spurious differences between group losses. We present CTC-DRO, which addresses the shortcomings of the group DRO objective by smoothing the group weight update to prevent overemphasis on consistently high-loss groups, while using input length-matched batching to mitigate CTC’s scaling issues. We evaluate CTC-DRO on the task of multilingual automatic speech recognition (ASR) across five language sets from the ML-SUPERB 2.0 benchmark. CTC-DRO consistently outperforms group DRO and CTC-based baseline models, reducing the worst-language error by up to $47.1\%$ and the average error by up to $32.9\%$. CTC-DRO can be applied to ASR with minimal computational costs, and offers the potential for reducing group disparities in other domains with similar challenges.
@@ 1 Introduction
To address these issues, we present CTC-DRO, which optimizes a generalization of the group DRO objective, specifically by smoothing the up-weighting of high-loss groups. This new objective prevents overemphasis on groups with consistently and disproportionately high training losses. Also, we use length-matched group losses to mitigate the scaling properties of CTC. We evaluate CTC-DRO using language subsets randomly selected from the ML-SUPERB 2.0 (Shi et al., 2024) benchmark. In this setting, CTC-DRO models outperform both group DRO and CTC-based baseline models across five language sets. Specifically, CTC-DRO models reduce the error rate of the worst-performing language in all of the five sets, with improvements of up to $47.1\%$, while also improving the average performance across all languages by up to $32.9\%$. Our code and newly trained models are publicly available at https://github.com/Bartelds/ctc-dro.
@@ 5 Results
1 MMS Baseline 60.8 (nan) 23.4 97.4
group DRO 86.6 (nan) 30.5 78.7
CTC-DRO 56.8 (nan) 22.9 95.8
XLS-R Baseline 64.9 (cmn) 25.2 92.6
group DRO 78.4 (nan) 30.0 87.8
CTC-DRO 57.6 (nan) 22.5 89.5
2 MMS Baseline 49.4 (yue) 15.8 98.4
group DRO 55.5 (yue) 20.7 98.2
CTC-DRO 44.4 (yue) 15.0 96.2
XLS-R Baseline 68.8 (yue) 19.0 94.2
group DRO 58.8 (yue) 21.6 87.0
CTC-DRO 45.0 (yue) 15.8 89.3
3 MMS Baseline 34.2 (kor) 16.1 98.5
group DRO 34.0 (kor) 22.0 98.7
CTC-DRO 31.3 (khm) 15.3 98.7
XLS-R Baseline 33.2 (khm) 17.0 99.2
group DRO 38.0 (khm) 25.1 97.2
CTC-DRO 32.2 (khm) 17.7 97.9
4 MMS Baseline 24.0 (snd) 14.4 87.9
group DRO 21.8 (urd) 14.9 91.9
CTC-DRO 18.4 (urd) 12.9 87.3
@@
5 MMS Baseline 90.0 (jpn) 26.0 96.3
group DRO 62.2 (jpn) 29.2 67.0
CTC-DRO 57.5 (jpn) 24.3 90.5
XLS-R Baseline 114.8 (jpn) 29.9 89.0
group DRO 92.9 (jpn) 36.8 57.7
CTC-DRO 71.5 (jpn) 23.8 91.0
@@
For each language set, CTC-DRO models achieve a lower CER for the worst-performing language compared to the baseline and group DRO models. The largest improvement is obtained on set 2 using XLS-R using all available data, showing a relative CER reduction of $47.1\%$ compared to the baseline model. Note that CTC-DRO also results in the best average CER in 13 out of 14 settings (seven sets with two models each) compared to both the baseline and group DRO models, leading to relative CER reductions up to $32.9\%$. The only exception is for XLS-R in balanced set 3, where the average CER is slightly worse with CTC-DRO (17.7%) than the baseline (17.0%). In terms of LID accuracy, CTC-DRO models improve over the baseline models in seven out of 14 settings. In the remaining settings, the LID accuracy already exceeds 87%, leaving little room for further improvement with CTC-DRO.
In contrast, group DRO worsens the CER of the worst-performing language in seven out of 14 settings compared to the baseline model, with the highest relative CER increase of $57.5\%$ on set 2 using MMS trained on all available training data. Also, group DRO increases the average CER compared to the baseline in all settings. This finding shows the ineffectiveness of the original group DRO formulation in this challenging setting, and the substantial added robustness of the modifications in CTC-DRO.
In four settings, the worst-performing language changes between the baseline and CTC-DRO models. For example, in set 3 with MMS trained on balanced data, it shifts from Korean to Khmer. As shown in Table 7, the CTC-DRO model reduces the CER for Korean from 34.2 to 27.6, while the CER for Khmer remains unchanged at 31.3. Overall, CTC-DRO consistently improves the performance of the worst-performing language, occasionally at a small cost to the performance of other languages, while still achieving a lower CER on average (see Appendix B).
@@ 6.1 Ablation Study
We find that removing either component from CTC-DRO leads to a substantial decrease in performance (see Table 4, detailed results in Appendix B). Specifically, the CER of the worst-performing language increases by up to $171.6\%$ and the average CER by up to $302.9\%$ compared to a model trained using the complete CTC-DRO algorithm. The ablation study also shows that the smoothed maximization objective has the strongest effect on reducing both the CER of the worst-performing language and the average CER.
@@ Appendix B Results
1 MMS Baseline 8.4 52.4 60.8 3.6 13.3 1.8 23.4 97.4
group DRO 20.6 48.6 86.6 4.3 16.7 6.2 30.5 78.7
CTC-DRO 10.5 46.1 56.8 3.7 17.9 2.3 22.9 95.8
XLS-R Baseline 7.3 64.9 60.8 3.1 13.4 1.8 25.2 92.6
group DRO 27.4 48.9 78.4 3.7 14.9 6.6 30.0 87.8
CTC-DRO 7.8 50.7 57.6 3.0 14.2 1.8 22.5 89.5
@@
2 MMS Baseline 0.2 21.8 9.0 5.9 8.2 49.4 15.8 98.4
group DRO 11.8 29.7 10.8 6.2 10.2 55.5 20.7 98.2
CTC-DRO 0.5 22.1 8.8 5.5 8.6 44.4 15.0 96.2
XLS-R Baseline 0.1 20.6 10.9 4.6 8.9 68.8 19.0 94.2
group DRO 12.7 28.5 14.4 5.1 10.2 58.8 21.6 87.0
CTC-DRO 0.5 21.5 12.6 5.2 10.0 45.0 15.8 89.3
@@
3 MMS Baseline 31.3 12.2 34.2 7.4 2.5 9.0 16.1 98.5
group DRO 33.2 19.1 34.0 22.4 9.8 13.5 22.0 98.7
CTC-DRO 31.3 12.0 27.6 8.1 2.3 10.2 15.3 98.7
XLS-R Baseline 33.2 13.3 32.3 8.7 3.7 11.0 17.0 99.2
group DRO 38.0 23.9 35.5 26.6 11.9 14.9 25.1 97.2
CTC-DRO 32.2 14.8 31.9 10.1 5.0 12.0 17.7 97.9
@@
5 MMS Baseline 5.4 11.1 30.2 90.0 12.0 7.2 26.0 96.3
group DRO 27.6 27.0 32.6 62.2 17.6 8.4 29.2 67.0
CTC-DRO 10.9 15.4 39.2 57.5 13.2 9.3 24.3 90.5
XLS-R Baseline 4.8 9.2 33.2 114.8 10.5 7.1 29.9 89.0
group DRO 29.1 26.8 46.1 92.9 16.5 9.3 36.8 57.7
CTC-DRO 5.7 9.6 38.6 71.5 10.1 7.3 23.8 91.0